# Budgeted Online Continual Learning by Adaptive Layer Freezing and Frequency-based Sampling

### v1

@@ A.13 Comparison between aL and Other Freezing Methods
A.14 Applying Adaptive Layer Freezing to Attention-based Model
@@
A.15 Applying aL-SAR across Diverse Network Architectures
@@
A.16 Details about $A_{\text{AUC}}$
A.17 Detailed Information on the Calculation of $\mathcal{S}_{y_{i},y_{j}}$
We calculate the cosine similarity of the gradients between all pairs of samples in the training batch and update the EMA estimate of $S_{y_{i}y_{j}}$ using it. To further reduce the computational cost of calculating similarity, we use only $0.05\%$ of the model parameters for the calculation of similarity, since the gradient distribution of the subset of randomly selected weights is similar to the gradient of the entire weight set (Li et al., 2022).
A.18 Details about Measuring FLOPs
@@
A.19 Details about the Memory bBudget in Total-constrained CL
@@
A.20 Experimental Results in Memory Infinite Setup
@@
A.21 Comparison with Retrieval-Based CL Methods
(a) CIFAR-10 Gaussian task setup                          (b) CIFAR-10 Disjoint task setup
(a) CIFAR-100 Gaussian task setup                          (b) CIFAR-100 Disjoint task setup
A.22 Effect of Temperature $T$
A.23 Comparison with SOTA Efficient CL Methods
(a) CIFAR-10 Gaussian task setup                                      (b) CIFAR-10 Disjoint task setup
(a) CIFAR-100 Gaussian task setup                                      (b) CIFAR-100 Disjoint task setup
A.24 Toy Experiments on Similarity-Aware Retrieval
@@
A.25 Limitations and Future Work

### v2

@@ A.15 Comparison between aL and Other Freezing Methods
A.16 Applying Adaptive Layer Freezing to Attention-based Model
@@
A.17 Applying aL-SAR across Diverse Network Architectures
@@
A.18 Detailed Information on the Calculation of $\mathcal{S}_{y_{i},y_{j}}$
We calculate the cosine similarity of the gradients between all pairs of samples in the training batch and update the EMA estimate of class-wise gradients, which are then used for calculating $S_{y_{i}y_{j}}$. To further reduce the computational cost of calculating similarity, we use only $0.05\%$ of the model parameters for the calculation of similarity, since the gradient distribution of the subset of randomly selected weights is similar to the gradient of the entire weight set (Li et al., 2022).
Specifically, we randomly select 0.05% of the model parameters across all layers before training. During training, we use the gradients of the pre-selected, which are unfrozen by adaptive layer freezing, to update the class-wise gradients. Specifically, we employ an EMA estimation of the class-wise gradients, updating them every batch iteration using the gradients of selected parameters in an EMA manner. While not all EMA-estimated class-wise gradients are updated every iteration (since gradients of frozen layers are not computed), all elements of the estimated class-wise gradients are updated every $m$ iterations, as we unfreeze all layers every $m$ iterations, as mentioned in Sec. A.1. To form the similarity matrix, we first compute the element-wise product of class-wise gradients and then average them over all class-wise pairs.
A.19 Comparison with Retrieval-Based CL Methods
(a) CIFAR-10 Gaussian task setup                          (b) CIFAR-10 Disjoint task setup
(a) CIFAR-100 Gaussian task setup                          (b) CIFAR-100 Disjoint task setup
A.20 Effect of Temperature $T$
A.21 Details about Measuring FLOPs
@@
A.22 Experimental Results in Memory Infinite Setup
(a) CIFAR-10 Gaussian task setup                       (b) CIFAR-10 Disjoint task setup
(a) CIFAR-100 Gaussian task setup                       (b) CIFAR-100 Disjoint task setup
A.23 Details about the Memory bBudget in Total-constrained CL
@@
A.24 Details about $A_{\text{AUC}}$
A.25 Comparison of Forgetting
[…]
ER 61.58 114,014
ER-MIR 61.05
DER++ 60.21
LiDER 59.21
X-DER 59.72
MEMO 59.25
CAMA 60.02
CCL-DC 61.26
aL-SAR (Ours) 56.28 947,128
Table 22: Comparison of KLR and KGR on ImageNet-1K Gaussian scheduled setup.
We compare the forgetting of aL-SAR with other baselines on CIFAR-10, CIFAR-100, and ImageNet-1K, and summarize the results in Fig.18, and Tab. 22, respectively. Specifically, in disjoint setup, we report $F_{last}$ (Chaudhry et al., 2018), following  Bang et al. (2021); Koh et al. (2022). In the Gaussian setup, however, the continuous shift in data distribution lacks explicit task boundaries, making traditional metrics like $F_{last}$ (Chaudhry et al., 2018) unsuitable. Therefore, we report the Knowledge Loss Ratio (KLR) (Koh et al., 2023), which do not require task boundaries.
As shown in Fig.18, forgetting decreases as training FLOPs increase. We believe this is because the increased computational budget allows the model to sufficiently train on previously encountered data before being exposed to new data.
@@ A.25 Comparison of Forgetting
Figure 18: Comparison of forgetting on Gaussian and Disjoint CL setup in CIFAR-10 and CIFAR-100.
A.26 Toy Experiments on Similarity-Aware Retrieval
@@
A.27 Limitations and Future Work